A uniform beam 6.0m and weighing 40N at the center,rests on two supports, P and Q placed 1.0m from each end of the beam. Weights of 100N and 80N are hung from the ends of the beam near P and Q respectively. Calculate the reactions at the supports P and Q.


1
Expert's answer
2021-04-13T06:31:58-0400

"\text{Let A and B these are the ends of the beam and }"

"\text{point C is the middle of the beam, then}"

"AP = 1;PC=2;PQ=4;PB = 5;"

"\text{By the condition of the problem:}"

"\vec F_a= -100;\vec F_c=-40;\vec F_b=-80;"

"\text{let } V_p \text{ and }V_q\text{ support reactions}"

"\Sigma M_p=0;\Sigma M_q=0;"

"\Sigma M_p = - F_a * AP + F_c * PC + F_b* PB - V_q * PQ ;"

"\Sigma M_p= -100+40*2+80*5-4*V_q" ;

"V_q =95 N"

"\Sigma M_q=-F_a*AQ-F_c*PC+F_B*PB+V_p*PQ;"

"V_p=\frac{ ( - 100 * 5 - 40 * 2 + 80 * 1)} { - 4} = 125N"

Answer:"V_p=125 N;V_q = 95N"